# Locally Convex Global Loss Network for Decision-Focused Learning

### v2

@@ Abstract
In decision-making problem under uncertainty, predicting unknown parameters is often considered independent of the optimization part. Decision-focused Learning (DFL) is a task-oriented framework to integrate prediction and optimization by adapting predictive model to give better decision for the corresponding task. Here, an inevitable challenge arises when computing gradients of the optimal decision with respect to the parameters. Existing researches cope this issue by smoothly reforming surrogate optimization or construct surrogate loss function that mimic task loss. However, they are applied to restricted optimization domain. In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global surrogate loss model which can be implemented in a general DFL paradigm. LCGLN learns task loss via partial input convex neural network which is guaranteed to be convex for chosen inputs, while keeping the non-convex global structure for the other inputs. This enables LCGLN to admit general DFL through only a single surrogate loss without any sense for choosing appropriate parametric forms. We confirm effectiveness and flexibility of LCGLN by evaluating our proposed model with three stochastic decision-making problems.
@@ 1 Introduction
Figure 1: A model training pipeline for PFL, DFL and surrogate DFL. PFL trains predictive model by minimizing the prediction loss. DFL directly delivers gradients minimizing the task loss. Surrogate DFL first learns a surrogate loss model that follows the true task loss well by sampling predictions and its task losses. Then, it trains the predictive model conveying useful gradients in an end-to-end manner.
@@
Decision-making problems under uncertainty arise in various real-world applications such as production optimization, energy planning, and asset liability management (Shiina and Birge 2003; Carino et al. 1994; Fleten and Kristoffersen 2008; Delage and Ye 2010; Garlappi, Uppal, and Wang 2007). These problems involve two main tasks: Prediction and Optimization. The Prediction task aims to create a model for uncertainty and estimate unknown parameters from some input features. This task is often performed using various machine learning (ML) techniques such as regression or deep learning. In the Optimization task, the corresponding optimization problem is solved using diverse off-the-shelf solvers with the estimated parameters from the prediction task. For example, in asset liability management, asset returns are estimated in the prediction stage, and the optimal portfolio to repay the liability is obtained in the optimization stage based on these predictions.
@@
Prediction-focused learning (PFL) framework is the most commonly used approach to solve such problems, handling the two tasks in separate and independent steps. PFL first trains ML model on the input features to produce predictions that closely match the observed parameters. Subsequently, the decision-making problem is defined using the parameters predicted by the trained ML model and is solved to obtain an optimal decision. PFL is based on the underlying belief that accurate predictions will lead to good decisions. However, ML models frequently fail to achieve perfect accuracy, resulting in sub-optimal decisions. Consequently, in many applications, the prediction and optimization tasks are not distinct but are intricately linked and should ideally be considered jointly.
@@
Decision-focused learning (DFL) achieves the above purpose by directly training ML model to make predictions that lead to good decisions. DFL combines prediction and optimization tasks by creating an end-to-end system that aims to directly minimize task loss. Task loss is defined as the quality of decisions derived from the predictive model and therefore, depends on the optimal solution of the associated optimization problem. To train an ML model in this context, differentiating through the optimization problem is required to calculate the gradient of the task loss. This presents a key challenge in integrating prediction and optimization. However, differentiation may be impossible if the decision variable is discrete or if the objective function is discontinuous.
@@
While surrogate optimization still necessitates differentiation of the optimization problem, efforts have been made to develop solver-free surrogate loss functions that can effectively obtain gradients. We refer to this approach as Surrogate DFL. Recent successful works in Surrogate DFL include SPO (Elmachtoub and Grigas 2022), Contrastive Loss (Mulamba et al. 2020), LTR Loss (Mandi et al. 2022), LODL (Shah et al. 2022), EGL (Shah et al. 2024), LANCER (Zharmagambetov et al. 2024), and TaskMet (Bansal et al. 2024). SPO, Contrastive loss, and LTR loss are specifically designed for linear objectives while LODL can be implemented for general DFL problems. However, LODL creates and trains a surrogate with a specific parametric form for each data instance resulting in expensive computation, with decision quality heavily dependent on the chosen parametric surrogate form. While EGL and LANCER have partially addressed the computational cost issue by extending a local model to a global model, the challenge of selecting an appropriate parametric surrogate form still remains. Alongside the surrogates, TaskMet minimized the prediction error while maintaining the decision quality. We illustrated the overall procedures of PFL, DFL, and Surrogate DFL in Figure 1.
@@
In this paper, we propose Locally Convex Global Loss Network (LCGLN), which is a global and general surrogate loss for DFL. LCGLN adopts partial input convex neural network (PICNN) (Amos, Xu, and Kolter 2017) as a parametric surrogate form to approximate task loss. With PICNN, we guarantee the surrogate loss function to be convex for the chosen inputs, while maintaining the general structure for the others. Furthermore, users are not required to possess an artistic sense for selecting the suitable parametric function form for the task loss. Therefore, LCGLN can handle the general DFL problems using only one surrogate loss, regardless of the number of observed data. We demonstrate the capability of LCGLN on three stochastic decision-making problem, namely inventory stock problem, budget allocation problem and portfolio optimization problem and show that LCGLN learns task loss well with a single surrogate loss.
@@ 2 Related Works
Various methodologies based on gradient based learning in DFL have been developed (Mandi et al. 2023). Some approaches directly differentiate the constrained optimization problem to update the model parameters. For instance, stochastic optimization problems were tackled by directly differentiating convex QPs using KKT optimality conditions and employing Optnet (Amos and Kolter 2017) for efficient differentiation (Donti, Amos, and Kolter 2017). For general convex optimization problems, the differentiable solver Cvxpylayers (Agrawal et al. 2019) were developed.
@@
Recent works on DFL have focused on constructing differentiable surrogate loss models. Smart ”Predict, Then Optimize” (SPO) (Elmachtoub and Grigas 2022) proposed a convex SPO+ loss where the loss upper bounds the task loss. To update the predictive model, they obtained a subgradient of their proposed surrogate loss. NCE (Mulamba et al. 2020) used a noise contrastive approach (Gutmann and Hyvärinen 2010) to build a family of surrogate loss for linear objectives. LTR (Mandi et al. 2022) applied this approach on ranking problems. SO-EBM (Kong et al. 2022) used an energy-based differentiable layer to model the conditional probability of the decision. The energy-based layer plays as a surrogate of the optimization problem. LODL (Shah et al. 2022) used supervised machine learning to locally construct surrogate loss models to represent task loss. They first sampled predictions near true instances for each instance and built a surrogate loss model respectively. They proposed three families for the model (e.g. WeightedMSE, Quadratic) to design a parametric surrogate loss models. EGL (Shah et al. 2024) extends LODL to global surrogate model and LANCER (Zharmagambetov et al. 2024) learns such handcraft global surrogate loss via actor-critic framework.
@@ 3 Preliminaries
In this section, we motivate necessity of DFL approach from simple knapsack example. We also briefly summarize surrogate loss models and input convex neural networks.
Figure 2: A simple example of knapsack problem. There are two items valued \$40, \$30 each, marked with yellow star. We predict the value of items and choose the higher one. Blue dots and red crosses are predicted values representing good and bad decisions respectively. PFL gives same prediction loss for every prediction while DFL gives \$10 loss in red cross and 0 in blue.
@@
Prediction focused learning (PFL) and decision focused learning (DFL) are two big learning pipelines for the decision making problems under uncertainty as in Figure 1. Decision making problems can be considered in two big stages: model training stage and evaluation stage. We would like to focus on the model training stage. PFL learns the predictive model focusing on the prediction from the predictive model while DFL focuses on the objective (or task loss for minimization). Starting with PFL, the model training stage is again divided in two steps, prediction and optimization. In prediction step, PFL learns a predictive model by minimizing the prediction loss such as MSE. In optimization step, it solves a constrained optimization problem using the prediction of predictive model. By contrast, the predictive model in DFL is trained in an end-to-end manner. It focuses on learning the model to make good decisions (or actions), and therefore it uses task loss (or objective for maximization problem), rather than concentrating on minimizing the prediction loss.
To give a motivation in usage of DFL, we introduce a motivating example with a simple knapsack problem in Figure 2. Let’s consider two items: item 1 with value \$40 and item 2 with \$30 (true value, yellow star). Assume we can only select one item. Our objective is to predict the value of items by only observing the item features and choose one with a higher value. If we decide to choose item 1 (the optimal decision in this case), we would be happier, earning \$40. Choosing item 2 is a sub-optimal decision, leading to a relative \$10 loss. Suppose we predicted the values as blue dots and red crosses. In PFL point of view, the prediction loss of our prediction loss is equal to \$20. Any prediction on the grey circle will give same prediction loss, which is not informative. In DFL standpoint, blue dots return zero task loss while red crosses gives task loss of \$10. We show that minimizing the prediction loss may not lead to minimizing the task loss.
@@
DFL compute task loss gradients with respect to the input parameters by for example, directly differentiating the optimization problem (or optimizer) (Donti, Amos, and Kolter 2017). However, the optimizer cannot be differentiated easily in most of the general cases. To tackle this, methods using surrogate loss functions were proposed (Elmachtoub and Grigas 2022). Surrogate loss models are differentiable and it approximates the mapping between the prediction and the task loss. As a consequence, it can be used to calculate the gradient to update the predictive model efficiently.
@@
To update the predictive model via gradient descent w.r.t. $\theta$, one must find the gradient $\partial\mathcal{R}/\partial\theta$. Using the chain rule, the gradient can be decomposed to:
@@
Amos et al. (Amos, Xu, and Kolter 2017) proposed input convex neural network (PICNN) which ensures convexity with respect to the chosen inputs. PICNN is constructed by introducing additional weights to connect input layer to each hidden layer. Given this structure, a non-decreasing convex activation functions such as softplus and non-negativity of weights connecting between hidden layers are required to guarantee convexity of PICNN. LCGLN uses PICNN to richly represent the non-convex nature of the true loss mapping, and simultaneously ensure local convexity.
@@ 4 Locally Convex Global Loss Network
In various optimization problems, each has its own specific objectives known as task losses. Our goal is to devise a surrogate loss function capable of accurately representing the true task loss across a range of optimization problems. In this paper, we introduce LCGLN, where a single loss representation can replace the true task loss. While DFL methods suffer from differentiating through the optimization solver when the optimization problem is not smooth, our LCGLN is easily differentiable, and therefore can be easily back-propagated when updating the predictive model via gradients.
4.1 End-to-end Training Procedure using LCGLN
@@ 3. Train Predictive Model $\mathcal{M}_{\theta}$:
Here, we introduce the end-to-end training procedure for LCGLN. The goal is to learn a regret mapping $(\hat{y},y)\rightarrow\mathcal{R}(\hat{y},y)$ to obtain gradient $\partial\mathcal{R}/\partial\theta$ for updating the predictive model $\mathcal{M}_{\theta}$.
@@
To train the LCGLN in a supervised learning manner, we generate $K$ samples for each $N$ instances, i.e. for each instance $y_{i}$, we sample $\tilde{y}_{i}^{(1)},\tilde{y}_{i}^{(2)},...,\tilde{y}_{i}^{(K)}$.
Previous research has assumed that predictions $\tilde{y}$ will closely approximate the true labels $y$, typically generating sample predictions by simply adding Gaussian noise. To capture realistic predictions, we adopt the model-based sampling (MBS) approach (Shah et al. 2024). This involves constructing a sampling model $M_{\xi}$ that mirrors the architecture of the predictive model and training within a PFL paradigm using MSE. During the training, we take inferences to the intermediate model and generate samples.
Learning Global Surrogate Loss LCGLN
Our objective is to learn a mapping $(\hat{y},y)\rightarrow\mathcal{R}$. For the loss model to convey informative gradients, we claim that $\mathcal{L}_{\psi}$ should be:
  • Expressive: Good surrogate loss model should have sufficient number of parameters to accurately approximate the true task loss. A lack of expressiveness may highly lead to under-performance. We leverage neural networks to ensure sufficient expressiveness.
  • Easily Differentiable: We need to differentiate our loss model with respect to $\hat{y}$ for gradients training the predictive model. Discontinuous or non-differentiable loss models will highly likely to suffer in such task. We wish our model to be easily differentiable.
  • Locally Convex, but not Globally: We want the model to capture the highly non-convex structure of the true underlying loss mapping. At the same time, we claim the loss model should be locally convex for every instance $y_{i}$, where the local minima is formed when $\hat{y}=y$. We induce the local minima by adding $\{(y_{i},y_{i},0)\}_{i=1}^{N}$ to the dataset for every instance.
Table 1: The table contains normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on three stochastic optimization problems. The metric is lower the better with 0 meaning optimal. We bold-lettered the best performing results for each problem. We evaluate PFL, DFL, local and global surrogate loss models. We use 32 samples for the surrogate loss models. The global surrogate loss LCGLN outperformed the baselines across all three problems.
@@
Gathering above approaches, we finally propose our Locally Convex Global Loss Network (LCGLN), a neural network that optimizes over the convex inputs to the network and outputs the regret. LCGLN uses partial input convex neural network (Amos, Xu, and Kolter 2017) to learn surrogate loss convex to the designated inputs with sufficient number of parameters. For LCGLN, we do not need to choose the parametric loss form to be used; we can simply use a single LCGLN to replace task loss. Furthermore, we can easily differentiate the model by using a built-in backward functions and obtaining gradients. Most importantly, by constructing a surrogate loss convex in each instance $y_{i}$ while capturing the non-convex nature among instances, LCGLN can richly represent the true task loss.
Training Predictive Model $M_{\theta}$
The predictive model learns a mapping $x\rightarrow y$. We train our predictive model $M_{\theta}$ via gradient descent using the gradients provided by the global surrogate loss $\mathcal{L}_{\psi}$. The procedure from generating dataset to training the predictive model is summarized in Algorithm 1.

### v4

@@ Abstract
In decision-making problems under uncertainty, predicting unknown parameters is often considered independent of the optimization part. Decision-focused learning (DFL) is a task-oriented framework that integrates prediction and optimization by adapting the predictive model to give better decisions for the corresponding task. Here, an inevitable challenge arises when computing the gradients of the optimal decision with respect to the parameters. Existing research copes with this issue by smoothly reforming surrogate optimization or constructing surrogate loss functions that mimic task loss. However, they are applied to restricted optimization domains. In this paper, we propose Locally Convex Global Loss Network (LCGLN), a global surrogate loss model that can be implemented in a general DFL paradigm. LCGLN learns task loss via a partial input convex neural network which is guaranteed to be convex for chosen inputs while keeping the non-convex global structure for the other inputs. This enables LCGLN to admit general DFL through only a single surrogate loss without any sense for choosing appropriate parametric forms. We confirm the effectiveness and flexibility of LCGLN by evaluating our proposed model with three stochastic decision-making problems.
@@ 1 Introduction
Figure 1: A model training pipeline for PFL, DFL, and surrogate DFL. PFL trains the predictive model by minimizing the prediction loss. DFL directly delivers gradients minimizing the task loss. Surrogate DFL first learns a surrogate loss model that follows the true task loss by sampling predictions and its task losses. Then, it trains the predictive model to convey useful gradients derived from the trained surrogate loss model in an end-to-end manner.
@@
Decision-making problems under uncertainty arise in various real-world applications such as production optimization, energy planning, and asset liability management (Shiina and Birge 2003; Carino et al. 1994; Fleten and Kristoffersen 2008; Delage and Ye 2010; Garlappi, Uppal, and Wang 2007). These problems involve two main tasks: Prediction and Optimization. The Prediction task aims to create a model for uncertainty and estimate unknown parameters from some input features. This task is often performed using machine learning (ML) techniques such as regression or deep learning. In the Optimization task, the corresponding optimization problem is solved using diverse off-the-shelf solvers with the estimated parameters from the prediction task. For example, in asset liability management, asset returns are estimated in the prediction stage, and the optimal portfolio to repay the liability is obtained in the optimization stage based on these predictions.
@@
Prediction-focused learning (PFL) framework is the most commonly used approach to solve such problems, handling the two tasks in separate and independent steps. PFL first trains the ML model on the input features to produce predictions that closely match the observed parameters. Subsequently, the decision-making problem is defined using the parameters predicted by the trained ML model and is solved to obtain an optimal decision. PFL is based on the underlying belief that accurate predictions will lead to good decisions. However, ML models frequently fail to achieve perfect accuracy, resulting in sub-optimal decisions. Consequently, in many applications, the prediction and optimization tasks are not distinct but are intricately linked and should ideally be considered jointly.
@@
Decision-focused learning (DFL) achieves the above purpose by directly training ML models to make predictions that lead to good decisions. DFL combines prediction and optimization tasks by creating an end-to-end system that aims to minimize task loss directly. Task loss is defined as the quality of decisions derived from the predictive model and therefore, depends on the optimal solution of the associated optimization problem. To train an ML model in this context, differentiating through the optimization problem is required to calculate the gradient of the task loss. This presents a key challenge in integrating prediction and optimization. However, differentiation may be impossible if the decision variable is discrete or the objective function is discontinuous.
@@
While surrogate optimization still necessitates differentiation of the optimization problem, efforts have been made to develop solver-free surrogate loss functions that can effectively obtain gradients. We refer to this approach as Surrogate DFL. Recent successful works in Surrogate DFL include SPO (Elmachtoub and Grigas 2022), Contrastive Loss (Mulamba et al. 2020), LTR Loss (Mandi et al. 2022), LODL (Shah et al. 2022), EGL (Shah et al. 2024), LANCER (Zharmagambetov et al. 2024), and TaskMet (Bansal et al. 2024). SPO, Contrastive loss, and LTR loss are specifically designed for linear objectives while LODL can be implemented for general DFL problems. However, LODL creates and trains a surrogate with a specific parametric form for each data instance resulting in expensive computation, with decision quality heavily dependent on the chosen parametric surrogate form. While EGL and LANCER have partially addressed the computational cost issue by extending a local model to a global model, the challenge of selecting an appropriate parametric surrogate form still remains. Alongside the surrogates, TaskMet minimized the prediction error while maintaining the decision quality. We illustrated the model training pipeline for PFL, DFL, and Surrogate DFL in Figure 1.
@@
In this paper, we propose Locally Convex Global Loss Network (LCGLN), which is a global and general surrogate loss for DFL. LCGLN adopts partial input convex neural network (PICNN) (Amos, Xu, and Kolter 2017) as a parametric surrogate form to approximate task loss. With PICNN, we guarantee the surrogate loss function to be convex for the chosen inputs while maintaining the general structure for the others. Furthermore, users are not required to possess an artistic sense for selecting the suitable parametric function form for the task loss. Therefore, LCGLN can handle the general DFL problems using only one surrogate loss, regardless of the number of observed data. In the experiment section, we demonstrate the capability of LCGLN on three stochastic decision-making problems, namely inventory stock problem, budget allocation problem, and portfolio optimization problem. We show that LCGLN learns task loss well with a single surrogate loss.
@@ 2 Related Works
Various methodologies based on gradient-based learning in DFL have been developed (Mandi et al. 2023). Some approaches directly differentiate the constrained optimization problem to update the model parameters. For instance, stochastic optimization problems were tackled by directly differentiating convex QPs using KKT optimality conditions and employing Optnet (Amos and Kolter 2017) for efficient differentiation (Donti, Amos, and Kolter 2017). For general convex optimization problems, the differentiable solver Cvxpylayers (Agrawal et al. 2019) were developed.
@@
Recent works on DFL have focused on constructing differentiable surrogate loss models. Smart ”Predict, Then Optimize” (SPO) (Elmachtoub and Grigas 2022) proposed a convex SPO+ loss where the loss upper bounds the task loss. To update the predictive model, they obtained a subgradient of their proposed surrogate loss. NCE (Mulamba et al. 2020) used a noise contrastive approach (Gutmann and Hyvärinen 2010) to build a family of surrogate loss for linear objectives. LTR (Mandi et al. 2022) applied this approach to ranking problems. SO-EBM (Kong et al. 2022) used an energy-based differentiable layer to model the conditional probability of the decision. The energy-based layer acts as a surrogate of the optimization problem. LODL (Shah et al. 2022) used supervised machine learning to locally construct surrogate loss models to represent task loss. They first sampled predictions near true instances for each instance and built a surrogate loss model respectively. They proposed three families for the model (e.g. WeightedMSE, Quadratic) to design parametric surrogate loss models. EGL (Shah et al. 2024) extends LODL to a global surrogate model and LANCER (Zharmagambetov et al. 2024) learns such handcraft global surrogate loss via actor-critic framework.
@@ 3 Preliminaries
In this section, we motivate the necessity of DFL approach from a simple knapsack example. We also briefly summarize surrogate loss models and input convex neural networks.
Figure 2: A simple example of a knapsack problem. There are two items valued \$40, \$30 each, marked with a yellow star. We predict the value of items and choose the higher one. Blue dots and red crosses are predicted values representing good and bad decisions respectively. PFL gives the same prediction loss for every prediction while DFL gives \$10 loss in red cross and 0 in blue.
@@
Prediction-focused learning (PFL) and decision-focused learning (DFL) are two big learning pipelines for decision-making problems under uncertainty as in Figure 1. While PFL learns the predictive model focusing on the prediction from the predictive model, DFL focuses on the objective of the downstream optimization problem, commonly referred to task loss. Starting with PFL, the model training step is divided into two stages, prediction and optimization. In the prediction stage, PFL learns a predictive model by minimizing the prediction loss such as MSE. In the optimization stage, it solves an optimization problem using the prediction of the predictive model as parameters. By contrast, the predictive model in DFL is trained in an end-to-end manner. It prioritizes learning the model to make good decisions (or actions) by minimizing the task loss, rather than optimizing the prediction loss.
We introduce a motivating example with a simple knapsack problem in Figure 2 to give motivation in using DFL. Let’s consider two items: item 1 with a value of \$40 and item 2 with \$30. Two axes in Figure 2 represent the value of each item. The true optimal value $(40,30)$ is marked by a yellow star. Assume we can only select one item. Our objective is to predict the value of items by only observing the item features and choosing one with a higher value. If we choose item 1 (the optimal decision), we would be satisfied, earning \$40. Conversely, choosing item 2 would be a sub-optimal decision, resulting in a relative loss of \$10. Suppose we predict the values as blue dots and red crosses. From a PFL point of view, the prediction loss is equal to \$20. Any prediction on the gray circle results in the same prediction loss, which is not informative as the blue dots are correct decisions, whereas the red crosses are not. From a DFL standpoint, blue dots return zero task loss while red crosses give task loss of \$10. This example shows that minimizing the prediction loss may not lead to lower task loss.
@@
DFL computes task loss gradients with respect to the input parameters by, for example, directly differentiating the optimization problem (or optimizer) (Donti, Amos, and Kolter 2017). However, the optimizer cannot be differentiated easily in most cases. To tackle this, methods using surrogate loss functions were proposed (Elmachtoub and Grigas 2022). Surrogate loss models are differentiable and it approximates the mapping between the prediction and the task loss. Consequently, it can be used to calculate the gradient to update the predictive model efficiently.
@@
To update the predictive model via gradient descent w.r.t. $\theta$, one must find the gradient $\partial\mathcal{R}/\partial\theta$. Using the chain rule, the gradient can be decomposed into:
@@
Amos et al. (Amos, Xu, and Kolter 2017) proposed an input convex neural network (PICNN) that ensures convexity with respect to the chosen inputs. PICNN is constructed by introducing additional weights to connect the input layer to each hidden layer. Given this structure, non-decreasing convex activation functions such as softplus and non-negativity of weights connecting between hidden layers are required to guarantee the convexity of PICNN. LCGLN uses PICNN to richly represent the non-convex nature of the true loss mapping, and simultaneously ensure local convexity.
@@ 4 Locally Convex Global Loss Network
In various optimization problems, each has its own specific objectives known as task losses. Our goal is to devise a surrogate loss function capable of accurately representing the true task loss across a range of optimization problems. In this paper, we introduce LCGLN, where a single loss representation can replace the true task loss. While DFL methods suffer from differentiating through the optimization solver when the optimization problem is not smooth, our LCGLN is easily differentiable and therefore can be readily backpropagated when updating the predictive model via gradients.
@@ 3. Train Predictive Model $\mathcal{M}_{\theta}$:
We now introduce the end-to-end training procedure for LCGLN. The goal is to learn a regret mapping $(\hat{y},y)\rightarrow\mathcal{R}(\hat{y},y)$ to obtain gradient $\partial\mathcal{R}/\partial\theta$ for updating the predictive model $\mathcal{M}_{\theta}$. The training consists of three main steps: generating samples, learning a global surrogate loss, and training a predictive model.
@@
To train the LCGLN in a supervised learning manner, we generate $K$ samples for each $N$ instance, i.e. for each instance $y_{i}$, we sample $\tilde{y}_{i}^{(1)},\tilde{y}_{i}^{(2)},...,\tilde{y}_{i}^{(K)}$.
Some previous research assumed that predictions $\tilde{y}$ would closely approximate the true labels $y$, typically generating sample predictions by simply adding Gaussian noise. However, the Gaussian sampling method may be challenging to apply for two reasons. First, in some cases, the true label $y$ is unknown. For example, in the inventory stock problem used in our experiments, the true $y$ represents the probability vector of each demand occurring. Yet, the observed data only provides a specific demand value realized according to that probability. In such cases, since the true $y$ is unknown, there is no target to which Gaussian noise can be added. Second, even if the true $y$ is known, determining the appropriate standard deviation for sampling around $y$ can be difficult, often requiring repeated tuning.
On account of this, we adopt the model-based sampling (MBS) approach (Shah et al. 2024). MBS involves constructing a sampling model $M_{\xi}$ that mirrors the architecture of the predictive model and training within a PFL paradigm using MSE. During the training, we take inferences from the intermediate model and generate samples. Unlike the Gaussian sampling, neither the true $y$ nor the standard deviation for the noise is required for MBS.
@@
Table 1: The table contains normalized test regret $\mathcal{R}_{test}/\mathcal{R}_{worst}$ with standard error mean (SEM) tested on three stochastic optimization problems. The metric is lower the better with an optimal regret of zero. The best-performing results for each problem are bold-lettered. We evaluate PFL, DFL, local, and global surrogate loss models. We use 32 samples for the surrogate loss models. The global surrogate loss LCGLN outperformed the baselines across all three problems.
Learning Global Surrogate Loss LCGLN
Our objective is to learn a mapping $(\hat{y},y)\rightarrow\mathcal{R}$ for conveying informative gradients of task loss. We propose leverage of partial input convex neural network (PICNN) (Amos, Xu, and Kolter 2017) as our Locally Convex Global Loss Network (LCGLN). Our motivation for using PICNN as a surrogate model for task loss is fourfold:
  • Expressiveness: A good surrogate loss model should have a sufficient number of parameters to accurately approximate the true task loss since a lack of expressiveness may lead to under-performance. Amos, Xu, and Kolter (2017) proved that a $k$-layer PICNN can represent any $k$-layer feedforward network. Since feedforward neural networks are known as universal approximators (Cybenko 1989; Funahashi 1989; Hornik 1991), this ensures the PICNN can accurately model task loss.
  • Easily Differentiable: We need to differentiate our loss model with respect to $\hat{y}$ for gradients training the predictive model. Discontinuous or non-differentiable loss models face significant challenges in such tasks. In contrast, PICNN can be easily differentiated by using built-in backward functions.
  • Locally Convex, but not Globally: We want the model to capture the highly non-convex structure of the true underlying task loss mapping. Task loss can be expressed as $f(x^{\ast}(\hat{y}),y)-f(x^{\ast}(y),y)$, where it achieves its minimum regret zero when $\hat{y}=y$. While task loss may not be convex in $\hat{y}$ for a given $y$ in general, we use PICNN to drive $\hat{y}$ towards the true value $y$ by eliminating local minima except $y$. This property enables the model to provide informative gradients for training the predictive model with a small sample size. Furthermore, we induce the local minima by adding $\{(y_{i},y_{i},0)\}_{i=1}^{N}$ to the dataset for every instance.
  • Generality: The adoption of PICNN helps generalize the function approximators and allows the approximation of differentiable optimization. Using PICNN, we can only focus on the overall network architecture and perform a simple hyperparameter search instead of requiring well-trained experts’ efforts in choosing the right parametric function forms for specific problems.
Training Predictive Model
The predictive model $M_{\theta}$ learns a mapping $x\rightarrow y$. To obtain gradients for the predictive model training, we first generated samples using the MBS approach. For each generated sample, we derived the true task loss in regret form and added the sample, instance, and regret pair to the dataset. Additionally, we included $\{(y_{i},y_{i},0)\}_{i=1}^{N}$ into the dataset to induce local minima for every instance $y_{i}$. Using this dataset, we trained the global surrogate loss model structured with PICNN.
Building on these steps, we now train the predictive model $M_{\theta}$ via gradient descent utilizing the gradients provided by the global surrogate loss $\mathcal{L}_{\psi}$. The entire procedure from generating the dataset to training the predictive model is summarized in Algorithm 1.